MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2016/12/05
OBSERVER: Alex Li
CONDITION: Windy & Cloudy
COMMENT: Dark images for 300s and 30s were taken. lmc2 was observed but with bad seeing (3.9) and cloudy image for lmc12
Compressor gas pressure: 1.75 MPa
Expected Observation Time: START 22:56 END 04:19


Go to the Seeing Log

Go to the Observation Log

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB-000
LMC13706-1370612600
SMC-000
TRANSIT-000
ALERT-000
GRB-000
TOTAL---12600 (00h10m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
00000000000


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
010000000001
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


TRANSIT FIELD STATISTICS

Field1Field2Field3
000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
23:00waitingwindy, cloudy
00:00waitingwindy, cloudy
01:00waitingwindy, cloudy
02:00observingobserve LMC, but cloudy later so no lmc12 & 14. Seeing (3.9) is bad for lmc2


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day